# Controllable seismic velocity synthesis using generative diffusion models

### v1

@@ 3 Numerical Experiments
The proposed method can handle a wide range of multi-modal conditions to control the velocity synthesis. Here, we report our tests on the OpenFWI Deng et al. (2021) dataset and showcase the flexibility of the proposed method to incorporate different conditions, specifically the class labels, well logs, and reflectivity images, to control the velocity synthesis. We utilize the 2D velocity models in these tests, which includes eight classes of models, so-called "FlatVel-A, "FlatVel-B", "CurveVel-A", "CurveVel-B", "FlatFault-A", "FlatFault-B", "CurveFault-A", "CurveFault-B", with samples shown in Figure 3. There are 336000 velocity samples in the dataset, and the model size of each velocity in our test is 64$\times$64 and the spatial interval is 10 m in horizontal and vertical directions. We train the diffusion model using an Adam optimizer with a learning rate of 1e-4. The batch size for the training is 1024, and the maximum training iteration is set to 200000. The backbone of the diffusion model is a U-Net with attention blocks Oktay et al. (2018). We use four feature map resolutions (from 64$\times$64 to 16$\times$8). The diffusion timesteps, $t$, is set to 1000 and embedded via a sinusoidal positional encoding into each residual block. The details of the backbone networks architecture are shown in Figure 4.
@@ 3.1 Unconditional Denoising Diffusion Probability models
In such a setting, the training process is focused on the denoising neural networks. We are storing the distribution $p(x)$ into the U-Net model. We evaluate the generation progress with 64 random Gaussian noise samples as input to the reverse progress, and Figure 5 shows the corresponding generated velocities. Without imposing any condition to affect the sampling process, the diffusion model produces velocities with the maximum diversity, which means the generated velocities can belong to any of the eight previously mentioned classes, or even be in a form that represents a combination of the classes, depending on their distribution.
@@ 3.2 Class-conditioned generation
In some scenarios, we would want to control the velocity generation based on prior information. This control can be useful in generating datasets for training as well as in the application of full waveform inversion. In this subsection, we first show the application of class-conditioned velocity synthesis. As mentioned earlier, there are eight classes in the velocity dataset. We incorporate this condition by using a classifier-free guidance diffusion model shown in Figure 6. For the class embedding, we use a small neural network with one hidden layer of size 256. As suggested in Ho and Salimans (2022), we jointly train the diffusion model with the condition imposed and without the condition by randomly setting the condition to zero with a probability of 0.5. The sampled results corresponding to their given classes are shown in Figure 6. We observe that the velocity synthesis can be controlled with the class labels, which is favorable to guide the diffusion model to generate the velocity we want. However, this also means the diversity of the velocity synthesis is limited within a given class.
@@
Another potential class option may involve the smoothness of the velocity model. This can be useful in full waveform inversion as we build the velocity model from smooth to sharp (high resolution) components. Thus, using the smoothness to control the velocity generation would be interesting. Here, to demonstrate the feasibility of the proposed method for such a scenario, we create three classes, from the original ”CurveFaultB” class, based on the level of smothness, including "CurveFaultB", "CurveFaultBSmooth5 (CFBSmooth5)", "CurveFaultBSmooth15 (CFBSmooth5)". "CFBSmooth5", for example, denotes the generated smoothed velocity by using a Gaussian filter applied to the original "CurveFaultB" model with a sigma of 5. The generated results are shown in Figure 7, and we can see that the proposed method can handle classifying such cases. Setting a target level of smoothness of the velocity model, the diffusion model generates a velocity with the right amount of smoothness. These generated smooth models correspond to the original distribution, which includes the smoother versions. In future work, we will incorporate such a diffusion model into the workflow of full waveform inversion.
@@ 3.3 Well-conditioned generation
Figure 8: The controllable samples with reconstruction-guided diffusion models, where the condition scale is 1.0 and we use 16 wells as conditions.
Figure 9: The controllable samples with reconstruction-guided diffusion models, where the condition scale is 8.0 and we use 16 wells as conditions.
However, when the number of well logs is reduced, we found that the generation is unstable. As shown in Figure 10, the generated velocities are not consistent with the provided well information. As denoted by the red box in Figure 10, there are small anomalies. This is caused by the gradient from the MSE loss between the well logs and generated velocities, which is insufficient to guide the sampling process to the target that satisfies the well logs as well as the prior distribution. In real-world applications, there are often only a few well-logs in the region, which means this approach may fail to produce an ideal velocity.
Figure 10: The controllable samples with reconstruction-guided diffusion models, where the condition scale is 32.0 and we use 4 wells as conditions.
The second one is the classifier-free generation. As mentioned earlier, this approach implicitly uses the gradient of the neural networks, whose inputs are the well logs, to guide the sampling process. However, it is more robust compared to the reconstruction-guided. In this approach, we can handle the case of a single well log, which is common in realistic problems in 2D surveys. As shown in Figure 11, with only one well to guide the sampling process, the generated velocity matches the well log at the well location. Compared to the class-guided velocity synthesis, this type of method provides a more feasible and more realistic scenario as we usually have one well in the region. To further test the diversity of the proposed method for this scenario, we show 16 samples generated by the same well log in Figure 12. As expected, the well imposes its information mainly at the location of the well, and away from the well, we are free to generate random models from the distribution. Thus, in the next section, we will show results corresponding to using the structural information to control the velocity synthesis.
Figure 11: The controllable samples with a classifier-free diffusion model, where we use one well as the condition.
Figure 12: The controllable samples with classifier-free diffusion models, where we use one well located at the horizontal location of 30 as a condition to generate 64 samples to demonstrate the diversity of the generation.
@@ 3.4 Seismic-image-conditioned generation
As mentioned earlier, spatial positional encoding is quite crucial for accurate velocity synthesis control. Here, we show in Figure 14 cases where the generator fails to adhere to the condition when spatial positional encoding is not used. We observed that although the generated velocities have similar structural information to the given reflectivity, they are not the same, and in some cases, they are out of symmetry with the reflectivity image. This is because the cross-attention can only capture the structural information without the proper positioning or direction. The spatial positional encoding mitigates this problem by injecting positional information on the reflectivity into the diffusion model, yielding more accurate controllable velocity synthesis.
Figure 13: The controllable samples conditioned by the corresponding reflectivity via cross-attention mechanism.
Figure 14: The controllable samples conditioned by the corresponding reflectivity via cross-attention mechanism without spatial positional encoding.
@@ 3.5 Integrated all-conditions for generation
Figure 15: The well logs and reflectivity used for controlling the velocity generation.
Figure 16: The controllable samples conditioned using class labels, well logs, and reflectivity images.
@@ 4 Discussion
An important question that may arise is how well these generatations work with out-of-distribution conditions. To test this, we focus our attention on the well-condition implementation shared in Section 3.3. In realistic scenarios, we may face the issue of out-of-distribution controllable velocity synthesis as the well velocities may exceed the minimum or maximum velocity of the training dataset. Here, we found that the proposed method can still handle such out-of-distribution cases, as shown in Figure 17. Although the generated velocities do not perfectly match the well due to its out-of-distribution nature, where the maximum velocity is 5250 $m/s$, and the minimum velocity is 460 $m/s$, it still provides a good approximation of the well velocity. Another out-of-distribution test is a continuously changing well velocity extracted from the field, and such rapid velocity changes are not represented in our trained models (see Figure 5). The well log is located at the North-Western Australia Continental shelf and is acquired by CGG, and the resulting generated velocity models are shown in Figure 18. We can see that the general velocity profiles from the generated results match the main trend of the well velocity. This demonstrates the potential application for realistic controllable velocity generation. Of course, if the diffusion model were trained on velocity models resembling this well, then the match would be stronger.
@@
Figure 17: The controllable samples with classifier-free diffusion model test using an out-of-distribution sample, with velocities out of the range of the training set. This well is located at the horizontal location of 30 as a condition to generate 64 samples to demonstrate the diversity of the generation.
@@
Figure 18: The controllable samples with a classifier-free diffusion model using another out-of-distribution condition from field data. The location of the well is at 25, and it corresponds to a 2D field marine dataset from the North-Western Australia Continental shelf acquired by CGG to generate 64 samples to demonstrate the diversity.
@@
The proposed method can handle multiple conditions in the velocity synthesis. However, in realistic applications, sometimes we may have conditions beyond the types studied in this paper. The proposed method, especially the use of cross-attention, inherently can handle multiple multi-modal conditions. We note that if the samples with conditions used for training are not sufficient, the proper way to incorporate the other conditions with limited samples (e.g., expert manual interpretation of the subsurface structures might be limited) is to use ControlNet Zhang et al. (2023). With this method, we can lock the pre-trained diffusion model and reuse the encoding layers as a backbone to learn diverse conditional controls, whose samples are not sufficient, to add more conditions.

### v2

@@ 3 Numerical Experiments
The proposed method can handle a wide range of multi-modal conditions to control the velocity synthesis. Here, we report our tests on the OpenFWI Deng et al. (2021) dataset and showcase the flexibility of the proposed method to incorporate different conditions, specifically the class labels, well logs, and reflectivity images, to control the velocity synthesis. We utilize the 2D velocity models in these tests, which include eight classes of models, so-called “FlatVel-A”, “FlatVel-B”, “CurveVel-A”, “CurveVel-B”, “FlatFault-A”, “FlatFault-B”, “CurveFault-A”, “CurveFault-B”, with samples shown in Figure 3. They have distinct features in structures and velocity variations with depth. “FlatVelA” and “CurveVelA” represent models without fault features, displaying a gradual increase in velocity with depth. In contrast, “FlatVelB” and “CurveVelB” also lack fault features, but here the velocity distribution is random along the depth profile. “FlatFaultA” and “CurveFaultA” introduce fault features with fewer discontinuities and more subtle variations in velocity. These are in contrast to “FlatFaultB” and “CurveFaultB”, where fault features are present with more discontinuities and more pronounced velocity changes. There are 336000 velocity samples in the dataset, and the model size of each velocity in our test is 64$\times$64 and the spatial interval is 10 m in horizontal and vertical directions. We train the diffusion model using an Adam optimizer with a learning rate of 1e-4. The batch size for the training is 1024, and the maximum training iteration is set to 200000. The backbone of the diffusion model is a U-Net with attention blocks Oktay et al. (2018). We use four feature map resolutions (from 64$\times$64 to 8$\times$8). The diffusion timesteps, $t$, is set to 1000 and embedded via a sinusoidal positional encoding into each residual block. The details of the backbone network architecture are shown in Figure 4. When applying cross-attention to incorporate the conditions, we replace the normalization and attention block in Figure 4 into a spatial transformer.
@@ 3.1 Unconditional Denoising Diffusion Probability models
In such a setting, the training process is focused on the denoising neural networks. We are using the U-Net model to learn the distribution $p(x)$. We evaluate the generation progress with 64 random Gaussian noise samples as input to the reverse progress, and Figure 5 shows the corresponding generated velocities. Without imposing any condition to affect the sampling process, the diffusion model produces velocities with the maximum diversity, which means the generated velocities can belong to any of the eight previously mentioned classes, or even be in a form that represents a combination of the classes, depending on their distribution.
@@ 3.2 Class-conditioned generation
In some scenarios, we would want to control the velocity generation based on prior information. This control can be useful in generating datasets for training as well as in the application of full waveform inversion. In this subsection, we first show the application of class-conditioned velocity synthesis. As mentioned earlier, there are eight classes in the velocity dataset. We incorporate this condition by using a classifier-free guidance diffusion model shown in Figure 2. For the class embedding, we use a small neural network with one hidden layer of size 256. As suggested in Ho and Salimans (2022), we jointly train the diffusion model with the condition imposed and without the condition by randomly setting the condition to zero with a probability of 0.5. The sampled results corresponding to their given classes are shown in Figure 6. We observe that the velocity synthesis can be controlled with the class labels, which is favorable to guide the diffusion model in generating the velocity we want. However, this also means the diversity of the velocity synthesis is limited within a given class.
@@
Another potential class option may involve the smoothness of the velocity model. This can be useful in full waveform inversion since we build the velocity model from smooth to sharp (high-resolution) components. Such a conditional generation provides the potential to control the smoothness of the generated velocity. It can help us incorporate the conditional diffusion models into diffusion regularized FWI, in which we can adjust the prior from the diffusion models to align with the model resolution (wavenumber) expected from FWI. Thus, using the smoothness to control the velocity generation would be interesting. Here, to demonstrate the feasibility of the proposed method for such a scenario, we create three classes, from the original ”CurveFaultB” class, based on the level of smothness, including "CurveFaultB", "CurveFaultBSmooth5 (CFBSmooth5)", "CurveFaultBSmooth15 (CFBSmooth5)". "CFBSmooth5", for example, denotes the generated smoothed velocity by using a Gaussian filter applied to the original "CurveFaultB" model with a sigma of 5. The generated results are shown in Figure 7, and we can see that the proposed method can handle classifying such cases. Setting a target level of smoothness of the velocity model, the diffusion model generates a velocity with the right amount of smoothness. These generated smooth models correspond to the original distribution, which includes the smoother versions. In future work, we will incorporate such a diffusion model into the workflow of full waveform inversion.
@@ 3.3 Well-conditioned generation
Figure 8: The controllable samples (right) with reconstruction-guided diffusion models, where the condition scale is 1.0 and we use 16 wells (left) as conditions.
Figure 9: The controllable samples (right) with reconstruction-guided diffusion models, where the condition scale is 8.0, and we use 16 wells (left) as conditions.
However, when the number of well logs is reduced, we found that the generation is unstable. As shown in Figure 10, although the generated velocities are still consistent with the given wells at the well location, they are poor in structure and horizontal continuity. As denoted by the red box in Figure 10, there are small anomalies. This is caused by the gradient from the reconstruction loss between the well logs and generated velocities, which is insufficient to guide the sampling process to the target that satisfies the well logs as well as the prior distribution. In real-world applications, there are often only a few well-logs in the region, which means this approach may fail to produce an ideal velocity.
Figure 10: The controllable samples (right) with reconstruction-guided diffusion models, where the condition scale is 32.0, and we use 4 wells (left) as conditions.
The second one is the classifier-free generation. As mentioned earlier, this approach implicitly uses the gradient of the neural networks, whose inputs are the well logs, to guide the sampling process. However, it is more robust compared to the reconstruction-guided. In this approach, we can handle the case of a single well log, which is common in realistic problems in 2D surveys. As shown in Figure 11, with only one well to guide the sampling process, the generated velocity matches the well log at the well location. Compared to the reconstruction-guided velocity synthesis, this type of method provides a more feasible and more realistic scenario as we usually have one well in the region. To further test the diversity of the proposed method for this scenario, we show 16 samples generated by the same well log in Figure 12. As expected, the well imposes its information mainly at the location of the well, and away from the well, we are free to generate random models from the distribution. Next, to test our method further on other types of conditions, we will show results corresponding to using the structural information to control the velocity synthesis.
Figure 11: The controllable samples (right) with a classifier-free diffusion model, where we use one well as the condition, and the profile comparison at the well location (left).
Figure 12: The controllable samples with classifier-free diffusion models, where we use one well located at the horizontal location of 30 as a condition to generate 16 samples (right) to demonstrate the diversity of the generation, and the profile comparison at the well location (left).
@@ 3.4 Seismic-image-conditioned generation
As mentioned earlier, spatial positional encoding is quite crucial for accurate velocity synthesis control. Here, we show in Figure 14 cases where the generator fails to adhere to the condition when spatial positional encoding is not used. We observed that although the generated velocities have similar structural information to the given reflectivity, they are not the same. In some cases, they are out of symmetry with the reflectivity image. This is because the cross-attention can only capture the structural information without the proper positioning or direction. The spatial positional encoding mitigates this problem by injecting positional information on the reflectivity into the diffusion model, yielding more accurate controllable velocity synthesis.
Figure 13: The controllable samples (right) conditioned by the corresponding reflectivity (left) via a cross-attention mechanism.
Figure 14: The controllable samples (right) conditioned by the corresponding reflectivity (left) via a cross-attention mechanism without spatial positional encoding.
@@ 3.5 Integrated all-conditions for generation
Figure 15: The well logs (left) and reflectivity (right) used for controlling the velocity generation.
Figure 16: The controllable samples (left) conditioned using class labels, well logs, and reflectivity images, and the profile comparisons (right) between the well log and generated the velocities at the well location.
@@ 4 Discussion
Another important question that may arise is how well these generatations work with out-of-distribution conditions. To test this, we focus our attention on the well-condition implementation shared in Section 3.3. In realistic scenarios, we may face the issue of out-of-distribution controllable velocity synthesis as the well velocities may exceed the minimum or maximum velocity of the training dataset. Here, we found that the proposed method can still handle such out-of-distribution cases, as shown in Figure 17. Although the generated velocities do not perfectly match the well due to its out-of-distribution nature, where the maximum velocity is 5250 $m/s$, and the minimum velocity is 460 $m/s$, it still provides a good approximation of the well velocity. To further demonstrate the performance of our conditional velocity synthesis, we use the Maximum Mean Discrepancy (MMD) (Smola et al., 2006) to quantify the difference between the original distributions and the distribution of the generated velocities. For the MMD calculation, we randomly sampled 1000 velocities from the original training dataset while generating 64 samples from out-of-distribution well logs. The MMD for this out-of-distribution test is 2.6454. With the same setting, the in-distribution test often results in a value less than 1.0. Because a large value of MMD denotes a larger distance between two distributions, the generated velocities are far away from the training distribution. Another out-of-distribution test is a continuously changing well velocity extracted from the field, and such rapid velocity changes are not represented in our trained models (see Figure 5). The well log is located at the North-Western Australia Continental shelf and is acquired by CGG, and the resulting generated velocity models are shown in Figure 18. We can see that the general velocity profiles from the generated results match the main trend of the well velocity, and the corresponding MMD is 1.4221. This demonstrates the potential application for realistic controllable velocity generation. Of course, if the diffusion model were trained on velocity models resembling this well, then the match would be stronger.
@@
Figure 17: The controllable samples (right) with classifier-free diffusion model test using an out-of-distribution sample, with velocities out of the range of the training set. This well (left) is located at the horizontal location of 30 as a condition to generate 16 samples to demonstrate the diversity of the generation.
@@
Figure 18: The controllable samples (right) with a classifier-free diffusion model using another out-of-distribution condition from field data. The location of the well (left) is at 25, and it corresponds to a 2D field marine dataset from the North-Western Australia Continental shelf acquired by CGG to generate 16 samples to demonstrate the diversity.
With respect to image-conditioned generation, as mentioned earlier, we can alternatively use the imaging results from applying reverse time migration (RTM) using a smooth version of the velocity models (possibly during training) or those obtained from migration velocity analysis. Thus, we extracted the 2000 velocity patches from the 3D Overthrust model, which has a reasonably complex geology and whose velocity varies spatially at the pixel level, to test the proposed method For such a limited number of velocity samples (2000) and limited diversity (extracted from one model), the resulting trained diffusion model is expected to suffer from poor generalization. The purpose here is to demonstrate the proposed method in controllable velocity generation using well logs and RTM images. The size of each velocity patch is 96$\times$96 with an interval of 25 $m$ in both vertical and horizontal directions. Though these are patches extracted from a 3D velocity model, we assume they represent individual velocity models for the same 2D area, as we aim to train the diffusion model to store their distribution. We smooth each velocity model with a Gaussian filter and then apply RTM to obtain the corresponding seismic imaging results, utilizing a 24Hz ricker wavelet and 10 sources, followed by a Laplacian filter. We use the same training configuration as that for the OpenFWI test but change the input resolution to 96$\times$96 and include the RTM images as well as random well logs as conditions. Given unseen well logs and RTM images from the test set (Figure 19), our proposed method can synthesize velocity models that align with the RTM images and correlate well with the well log, as demonstrated in Figure 20. This simple example demonstrates the effectiveness of the proposed method in such a scenario.
Figure 19: The well logs (left) and RTM results (right) used for controlling the velocity generation.
Figure 20: The generated samples (left) conditioned using well logs and RTM results, and the profile comparisons (right) between the well log and the generated velocities at the well location.
@@
Overall, the proposed method can handle multiple conditions in the velocity synthesis. However, in realistic applications, sometimes we may have conditions beyond the types studied in this paper. The proposed method, especially the use of cross-attention, can handle multiple multi-modal conditions inherently. We note that if the samples with conditions used for training are not sufficient, the proper way to incorporate the other conditions with limited samples (e.g., expert manual interpretation of the subsurface structures might be limited) is to use ControlNet Zhang et al. (2023). With this method, we can lock the pre-trained diffusion model and reuse the encoding layers as a backbone to learn diverse conditional controls, whose samples are not sufficient, to add more conditions.